# Automatically Generating UI Code from Screenshot: A Divide-and-Conquer-Based Approach

### v1

@@ Abstract.
In this paper, we propose DCGen, a divide-and-conquer-based approach to automate the translation of webpage design to UI code. DCGen starts by dividing screenshots into manageable segments, generating descriptions for each segment, and then reassembling them into complete UI code for the entire screenshot. We conduct extensive testing with a dataset comprised of real-world websites and various MLLMs and demonstrate that DCGen achieves up to a 14% improvement in visual similarity over competing methods. To the best of our knowledge, DCGen is the first segment-aware prompt-based approach for generating UI code directly from screenshots.
copyright: acmcopyrightjournalyear: 2024doi: XX.XXXX/XXXXXXX.XXXXXXXconference: Arxiv; preprint; 2024price: XX.XXisbn: 978-1-4503-XXXX-X/XX/XXsubmissionid: 1077copyright: none
@@ 1. Introduction
Our study curated a dataset comprising 1,000 top-ranked real-world websites. We evaluate the effectiveness of our methodology across several cutting-edge MLLMs, using 120 webpages sampled from the dataset. DCGen achieved up to 14% improvement in visual similarity between the original and generated websites compared to other design-to-code methods. We further demonstrate that DCGen is robust against various webpage complexities. Furthermore, DCGen is shown to generalize well across different MLLMs, demonstrating the effectiveness of the divide-and-conquer methodology.
@@
  • We initiated a motivating study to uncover types of errors in the LLM-powered design-to-code process, revealing the importance of visual segments in generation quality.
  • We propose DCGen with a divide-and-conquer approach. It starts with generating descriptions for individual image segments, then merges these solutions to produce the complete UI code.
  • The experiments on real-world webpages demonstrate the superiority of DCGen over other methods in both visual similarity and code similarity.
  • We release all datasets and code implementation for future research in this field.
@@ 2. Background
2.1. Concepts in Front-End Development
Front-end development concentrates on creating the user interface and enhancing the user experience of websites. It involves Hypertext Markup Language (HTML) for content structuring, Cascading Style Sheets (CSS) for styling, and optionally JavaScript for interactivity and dynamic webpage features.
2.1.1. HTML Elements & Tags
HTML is a markup language used in front-end web development to structure webpages. Tags in HTML define the document’s structure, like <h1> to <h6> for different heading levels, and <p> for paragraphs. Each HTML element includes an opening tag, content, and a closing tag, forming the basic block of a webpage. Below is a simplified HTML snippet containing six different elements:
<html>
<head>
<title>Sample Page</title>
</head>
<body>
<h1>Welcome to My Website</h1>
<p>This is a paragraph of text.</p>
</body>
</html>
2.1.2. CSS
CSS is used in front-end web development to style webpages and enhance the appearance of HTML elements. It allows developers to define unique styles for elements through selectors, properties, and values. For example, to change the color and font size of all <h1> elements, one developer might write:
h1 {
color: blue;
font-size: 24px;
2.2. Problem Definition
@@ 3. Dataset Collection
To avoid data leakage, we select the top 12 websites for the motivating study and use the remaining ones for evaluation.
@@ 4.1. Study Preparation
This step identifies the bounding boxes for each element (tag) in the code within the webpage screenshot, referred to as visual elements. Using Selenium WebDriver [^3] https://selenium-python.readthedocs.io/, we extract the bounding box coordinates for all elements, then merge bounding boxes located in similar positions. We also eliminate any bounding boxes that intersect, as webpage elements typically do not overlap. This collection of bounding boxes enables us to locate all HTML elements in a website’s screenshot. Overall, these webpages contain a total of 1,699 HTML elements, with individual webpages featuring between 4 and 434 elements, averaging 142 elements per site.
@@ 4.2. What failures does MLLM produce when converting design to code?
We ask annotators to assess the accuracy of visual elements in MLLM-generated webpages compared to the original ones. Figure 1 illustrates the annotation interface. The annotators are provided with pairs of screenshots, with the original on the left and the MLLM-generated version on the right. A red bounding box highlights the visual element in the original screenshot. Annotators are asked with two tasks: (1) Drawing the bounding box around the corresponding element in the MLLM-generated screenshot. (2) Describing any differences between the generated visual element and the original in short phrases.
@@
Result Analysis. After inspection, we discover that all the differences between the generated element and the original element can be characterized into three categories, that is, element omission, element distortion, and element misarrangement, and one element can have multiple errors at the same time. Figure 5 provides an example of each type of error, with the original website displayed above and the MLLM-generated version below.
@@
Figure 2. Examples of error cases (bottom).
@@ 4.3. Can we correct mistakes by visual segments?
The first case focuses on omitted elements. We isolate each missing element by cropping the screenshot to include only that specific element and then request the model to generate the corresponding code. Figure 3 depicts this process, where Figure 3(b) is the Frequently Asked Questions (FAQ) area generated from the full screenshot and Figure 3(c) is the FAQ area generated from the cropped screenshot of the original FAQ area. The results show that with these isolated inputs, the model successfully generates the previously omitted content. This demonstrates that the model is intrinsically capable of producing the required outputs but struggles with complex inputs.
@@
Refer to caption (b) Generated yellow dashed line area when input is the full screenshot. Refer to caption (c) Generated yellow dashed line area when input is the area’s screenshot.
(a) Input. The full image is the input of (b) and the yellow dashed line area is the input of (c).
Figure 3. An example of the experiment in the motivating study.
@@ 5. Methodology
(a) Original image
(b) Depth = 1
(c) Depth = 2
Figure 5. Examples of image subdivision of different depths.
@@ 5.1.1. Separation Line Detection
DCGen starts with identifying all separation lines (either vertical or horizontal) in a full screenshot. These separation lines must divide the screenshot into rectangular and meaningful pieces, where each of the pieces contains complete elements; otherwise, it would be more difficult to combine code for fractured pieces later. Traditional image segmentation techniques concentrate on distinguishing irregularly shaped objects from the background (Minaee et al., 2021), becoming unsuitable for dividing webpages.
@@
8:     diff_top $\leftarrow$ mean_abs_diff(upper, window[0])
9:     is_border_top $\leftarrow$ (diff_top ¿ diff_thr and portion(diff_top) ¿ portion_thr)
@@
11:     is_border_bottom $\leftarrow$ (diff_bottom ¿ diff_thr and portion(diff_bottom) ¿ portion_thr)
@@
(a) Explicit separation line between two sections.
(b) Implicit separation line between text contents.
Figure 6. Separation lines (marked with red dot line).
@@ 5.1.2. Screenshot Subdivision
We divide the screenshot into smaller segments by alternately and recursively applying horizontal and vertical divisions. Specifically, for each input screenshot, DCGen first attempts to divide the screenshot along horizontal separation lines. If no such lines are detected, it then proceeds to segment the screenshot using vertical separation lines by calling the separation line detector again. This process is applied recursively to each resulting segment until no further separation lines can be identified, or until a user-defined maximum recursion depth is reached. Figure 7 illustrates this subdivision procedure.
@@ 5.2. Assembly
As shown in Figure 4, for each leaf image segment, DCGen provides the MLLM (i.e., Leaf-solver MLLM) a website screenshot with a red rectangular bounding box specifying the part of the image it should focus on, then ask it to describe the image segment. For each parent image segment, DCGen provides the MLLM (i.e., Assembly MLLM) a screenshot with a bounding box, as well as all the descriptions generated from its children’s image segments. In the last step (merging for root node), DCGen calls the Generation MLLM to produce the complete UI code for full screen, based on the descriptions from its children segments.
Figure 8 illustrates the prompt details for the leaf-solver MLLM, assembly MLLM, and generation MLLM, respectively. We adapt the basic instructions based on the previous study by Si et al. (Si et al., 2024).
Figure 8. Prompt details for the leaf-solver MLLM, assembly MLLM, and generation MLLM, respectively.
@@ 6. Experiment
  • RQ1: How does DCGen perform in design-to-code generation compared to various direct prompting strategies?
  • RQ2: Which language modality (e.g., programming language or natural language) maximizes the performance of DCGen?
  • RQ3: How does the complexity of real-world websites impact the performance of DCGen?
  • RQ4: Can DCGen be generalized to other state-of-the-art (SOTA) MLLMs?
@@ 6.1.1. Backbone Models
To assess the performance of DCGen, we employ it to evaluate three SOTA MLLMs: GPT-4o (OpenAI, 2024a), Gemini-1.5 (Google, 2024), Claude-3 (Anthropic, 2024a). Access to all models is facilitated through their respective official APIs, the specific model numbers are 20240513 for GPT-4o, 20240229 for Claude-3-Sonnet, and Gemini-1.5-flash-latest accessed during May 2024. We randomly sampled 120 websites from our dataset to form the testing set. Table 3 shows the statistics of the testing dataset, the data is rich in HTML tags and diverse in overall complexity.
@@ 6.1.2. Parameter Setting
For MLLM model configurations, we set the temperature to 1, and the $max\_tokens$ parameter to the maximum allowable value for each model (i.e., 4096 for GPT-4o and Claude3-Sonnet, 8192 for Gemini-1.5-Flash), while other parameters are maintained at their default settings as specified in the corresponding API documentation (Google AI, 2024; OpenAI, 2024b; Anthropic, 2024b). For parameters in the image segmentation algorithm, we set var_thr=60, diff_thresh=5, diff_portion=0.3, and set the division depth to one.
@@ 6.1.3. Efficiency
Our experiments were conducted on a Windows 10 PC with an IntelCore i7-10870H CPU. We assess the performance of DCGen by first evaluating the speed of the Image subdivision algorithm, which completes the subdivision of a 1920x1080 pixel image in less than one second when the division depth is set to one. The response times for API requests across three models range from 5 to 15 seconds depending on the request length. On average, DCGen generates code for an entire webpage (1,370 tags) within 12 minutes, significantly faster than a person typing the tags manually (Typing.com, 2023). The leaf-solver MLLMs could work in parallel for (10x) faster computation, theoretically. This demonstrates DCGen’s practical efficiency in accelerating web development.
Figure 9. Propmt for baseline approaches.
@@ 6.1.4. Baselines
Figure 9 illustrates the details of prompting strategies used in our experiment as baselines. Chain-of-Thought (CoT) prompting (Wei et al., 2022a) generates a chain of thought for each question and then generates the corresponding code. For CoT, we use the ”let’s think step by step” instruction from Chae et. al (Chae et al., 2024). Self-refine prompting (Chen et al., 2023) let the model refine its own generated code. We adopt the self-refine prompting and direct promoting method from Si et al. (Si et al., 2024). We do not select previous deep-learning-based approaches as baselines since they are either tailored to the mobile app environment (Chen et al., 2018; Nguyen and Csallner, 2015; Moran et al., 2018a), not opensourced (Xu et al., 2021), do not take website screenshot as input or do not produce complete website code (Asiroglu et al., 2019; Beltramelli, 2018).
@@ 6.2.1. High-Level Similarity
For code similarity, we adopt the Bilingual Evaluation Understudy (BLEU) score (Papineni et al., 2002) to evaluate the closeness of generated code $C_{g}$ to the original code $C_{0}$. BLEU score is an automatic evaluation metric widely used in machine translation studies. It calculates the similarity of machine-generated translations and human-created references (i.e., ground truth). It is also widely used in code generation tasks to measure the similarity between generated code and the ground truth  (Chen et al., 2021; Yin et al., 2018; Wei et al., 2019; Ren et al., 2020; Lu et al., 2021). BLEU score is defined as the product of n-gram precision and a brevity penalty $BP$: $\text{BLEU}=BP\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right)$, where $p_{n}$ is the precision of the n-grams, $w_{n}$ is the weight (typically $\frac{1}{N}$), and $N$ is usually set to 4 (Sutskever et al., 2014). $BP$ is 1 if the candidate translation length $c$ exceeds the reference length $r$, and $\exp(1-\frac{r}{c})$ otherwise. BLEU scores range from 0 to 1, with higher scores indicating closer similarity to the reference translations.
For image similarity, we measure image similarity between $I_{0}$ and $I_{g}$ via CLIP Score  (Si et al., 2024). CLIP score relies on the similarity of their CLIP (Radford et al., 2021) embeddings, denoted as CLIP($I_{0}$, $I_{g}$). Specifically, we extract features using CLIP-ViT-B/32 after resizing the screenshots to squares. To eliminate noise introduced by text in the screenshots, we mask all detected text boxes using their bounding box coordinates with an inpainting algorithm implemented in OpenCV [^4] https://docs.opencv.org/4.3.0/df/d3d/tutorial_py_inpainting.html.
@@ 6.2.2. Fine-Grained Measurements
High-level metrics capture only the overall similarity of images and code, lacking the detail necessary to analyze specific model performances. To address this limitation, we adopt a suite of element-matching metrics for evaluating generated webpages in terms of text content, position, and color introduced by Si et al. (Si et al., 2024). Given reference and generated webpage screenshots $I_{0}$ and $I_{g}$, the algorithm detects visual element blocks and matches them using the Jonker-Volgenant algorithm. It then evaluates the similarity of the matching blocks across several aspects: (1) Block-match measures the ratio of matched block sizes to all block sizes, penalizing missing or hallucinated blocks; (2) Text similarity calculates character overlap using the Sørensen-Dice coefficient; (3) Color similarity employs the CIEDE2000 formula to evaluate perceptual color differences; and (4) Position similarity assesses the alignment of block centers. We use this set of metrics as indicators for the three types of errors found in Section 4. Specifically, we use (1) for measuring element inclusion (omission), (2) and (3) for measuring element fidelity (distortion), and (4) for indicating element organization (misarrangement).

### v2

@@ Abstract.
In this paper, we propose DCGen, a divide-and-conquer-based approach to automate the translation of webpage design to UI code. DCGen starts by dividing screenshots into manageable segments, generating descriptions for each segment, and then reassembling them into complete UI code for the entire screenshot. We conduct extensive testing with a dataset comprised of real-world websites and various MLLMs and demonstrate that DCGen achieves up to a 14% improvement in visual similarity over competing methods. Human evaluations show that DCGen can help developers implement webpages significantly faster and more similar to the UI designs. To the best of our knowledge, DCGen is the first segment-aware MLLM-based approach for generating UI code directly from screenshots.
Multi-modal Large Language Model, Code Generation, User Interface, Web Development
isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Our study curated a dataset comprising 1,000 top-ranked real-world websites. We evaluate the effectiveness of our methodology across several cutting-edge MLLMs, using 120 webpages sampled from the dataset. DCGen achieved up to 14% improvement in visual similarity between the original and generated websites compared to other design-to-code methods. We further demonstrate that DCGen is robust against various webpage complexities. Furthermore, DCGen is shown to generalize well across different MLLMs, demonstrating the effectiveness of the divide-and-conquer methodology. Human evaluations show that DCGen can produce more visually similar target webpages than competing methods and can help developers implement webpages faster and more similar to the UI design images.
@@
  • We initiated a motivating study to uncover types of errors in the MLLM-powered design-to-code process, revealing the importance of visual segments in generation quality.
  • We propose DCGen with a divide-and-conquer approach. It starts with generating descriptions for individual image segments, and then merges these solutions to produce the complete UI code.
  • The experiments on real-world webpages demonstrate the superiority of DCGen over other methods in both visual similarity and code similarity. Human evaluations also confirm that DCGen can help developers implement webpages faster and more similar to the UI designs.
  • We implement DCGen into a user-friendly demo tool and release all datasets and code implementation for future research in this field.
@@ 2. Background
Concepts in Front-End Development
Front-end development concentrates on creating the user interface and enhancing the user experience of websites. It involves Hypertext Markup Language (HTML) for content structuring, Cascading Style Sheets (CSS) for styling, and optionally JavaScript for interactivity and dynamic webpage features. HTML is a markup language used in front-end web development to structure webpages. Tags in HTML define the document’s structure, like <h1> to <h6> for different heading levels, and <p> for paragraphs. Each HTML element includes an opening tag, content, and a closing tag, forming the basic block of a webpage. CSS is used in front-end web development to style webpages and enhance the appearance of HTML elements. It allows developers to define unique styles for elements through selectors, properties, and values.
Problem Definition
@@ 4.1. Study Preparation
This step identifies the bounding boxes for each element (tag) in the code within the webpage screenshot, referred to as visual elements. Using Selenium WebDriver [^3] https://selenium-python.readthedocs.io/, we extract the bounding box coordinates for all elements and eliminate any bounding boxes that intersect, as webpage elements typically do not overlap. This collection of bounding boxes enables us to locate all HTML elements in a website’s screenshot. Overall, these webpages contain a total of 1,699 HTML elements, with individual webpages featuring between 4 and 434 elements, averaging 142 elements per site.
@@ 4.2. What failures does MLLM produce when converting design to code?
We recruit two graduate students in our university as annotators to assess the accuracy of visual elements in MLLM-generated webpages compared to the original ones. Figure 1 illustrates the annotation interface. The annotators are provided with pairs of screenshots, with the original on the left and the MLLM-generated version on the right. A red bounding box highlights the visual element in the original screenshot. Annotators are asked with two tasks: (1) Draw the bounding box around the corresponding element in the MLLM-generated screenshot. (2) Describe any differences between the generated visual element and the original in short phrases.
@@
Result Analysis. After inspection, we discover that all the differences between the generated element and the original element can be characterized into three categories, that is, element omission, element distortion, and element misarrangement, and one element can have multiple errors at the same time. Figure 3 provides an example of each type of error, with the original website displayed above and the MLLM-generated version below in each subfigure.
@@
Figure 2. Examples of error cases (bottom of each subfigure).
(d) Input. The full image is the input of (b) and the yellow dashed line area is the input of (c).
(e) Generated yellow dashed line area when input is the full screenshot.
(f) Generated yellow dashed line area when input is the area’s screenshot.
Figure 3. An example of the experiment in the motivating study.
@@ 4.3. Can we correct mistakes by visual segments?
The first case focuses on omitted elements. We isolate each missing element by cropping the screenshot to include only that specific element and then request the model to generate the corresponding code. Figure 3 depicts this process, where Figure 3b is the Frequently Asked Questions (FAQ) area generated from the full screenshot and Figure 3c is the FAQ area generated from the cropped screenshot of the original FAQ area. The results show that with these isolated inputs, the model successfully generates the previously omitted content. This demonstrates that the model is intrinsically capable of producing the required outputs but struggles with complex inputs.
@@ 5.1.1. Separation Line Detection
DCGen starts with identifying all separation lines (either vertical or horizontal) in a full screenshot. These separation lines must divide the screenshot into rectangular and meaningful pieces, where each of the pieces contains elements that are as complete as possible; otherwise, it would be more difficult to combine code for fractured pieces later. Traditional image segmentation techniques concentrate on distinguishing irregularly shaped objects from the background (Minaee et al., 2021), becoming unsuitable for dividing webpages.
@@
8:     diff_top $\leftarrow$ abs_diff(upper, window[0])
9:     is_border_top $\leftarrow$ portion(diff_top ¿ diff_thr) ¿ portion_thr
@@
11:     is_border_bottom $\leftarrow$ portion(diff_bottom ¿ diff_thr) ¿ portion_thr
@@
(a) Original image
(b) Depth = 1
(c) Depth = 2
Figure 5. Examples of image subdivisions of different depths.
(d) Explicit.
(e) Implicit.
Figure 6. Separation lines (red dot lines).
@@ 5.1.2. Screenshot Subdivision
We divide the screenshot into smaller segments by alternately and recursively applying horizontal and vertical divisions to mirror real-world development practices (Chopper, 2024). Specifically, for each input screenshot, DCGen first attempts to divide the screenshot along horizontal separation lines. If no such lines are detected, it then proceeds to segment the screenshot using vertical separation lines by calling the separation line detector again. This process is applied recursively to each resulting segment until no further separation lines can be identified, or until a user-defined maximum recursion depth is reached. Figure 7 illustrates this subdivision procedure.
@@ 5.2. Assembly
As shown in Figure 4, for each leaf image segment, DCGen provides the MLLM (i.e., Leaf-solver MLLM) a website screenshot with a red rectangular bounding box specifying the part of the image it should focus on, then ask it to describe the image segment. For each parent image segment, DCGen provides the MLLM (i.e., Assembly MLLM) a screenshot with a bounding box, as well as all the descriptions generated from its children’s image segments. In the last step (merging for root node), DCGen calls the Generation MLLM to produce the complete UI code for full screen, based on the descriptions from its children segments and the entire screenshot.
We adapt the basic instructions based on the previous study by Si et al. (Si et al., 2024). All the prompt details for the leaf-solver MLLM, assembly MLLM, and generation MLLM are made available in our artifact.
@@ 6. Experiment
  • RQ1: How does DCGen compare to various prompting strategies in design-to-code?
  • RQ2: How robust is DCGen for complex designs?
  • RQ3: How generalizable is DCGen for other state-of-the-art (SOTA) MLLMs?
  • RQ4: What is human’s feedback on DCGen in real-world application development workflows?
@@ 6.1.1. Backbone Models
To assess the performance of DCGen, we employ it to evaluate three SOTA MLLMs: GPT-4o (OpenAI, 2024a), Gemini-1.5 (Google, 2024), Claude-3 (Anthropic, 2024a). Access to all models is facilitated through their respective official APIs, the specific model numbers are 20240513 for GPT-4o, 20240229 for Claude-3-Sonnet, and Gemini-1.5-flash-latest accessed during May 2024. Due to a limitation on computational resources, we randomly sampled 120 websites from our dataset to form the testing set. Table 3 shows the statistics of the testing dataset, the data is rich in HTML tags and diverse in overall complexity.
@@ 6.1.2. Parameter Setting
For MLLM model configurations, we set the temperature to 0, and the $max\_tokens$ parameter to the maximum allowable value for each model (i.e., 4096 for GPT-4o and Claude3-Sonnet, 8192 for Gemini-1.5-Flash), while other parameters are maintained at their default settings as specified in the corresponding API documentation (Google AI, 2024; OpenAI, 2024b; Anthropic, 2024b). For parameters in the image segmentation algorithm detailed in Section 5.1, we set var_thr=60, diff_thresh=5, diff_portion=0.3, and set the division depth to one.
@@ 6.1.3. Efficiency
Our experiments were conducted on a Windows 10 PC with an IntelCore i7-10870H CPU. We assess the performance of DCGen by first evaluating the speed of the Image subdivision algorithm, which completes the subdivision of a 1920x1080 pixel image in less than one second when the division depth is set to one. On average, DCGen generates code for an entire webpage (1,370 tags) within 4 minutes. This demonstrates DCGen’s practical efficiency in accelerating web development.
@@ 6.1.4. Baselines
Chain-of-Thought (CoT) prompting (Wei et al., 2022a) generates a chain of thought for each question and then generates the corresponding code. For CoT, we use the ”let’s think step by step” instruction from Chae et. al (Chae et al., 2024). Self-refine prompting (Chen et al., 2023) let the model refine its own generated code via multi-turn conversation. We adopt the self-refine prompting and direct promoting method from Si et al. (Si et al., 2024). We do not select previous deep-learning-based approaches as baselines since they are either tailored to the mobile app environment (Chen et al., 2018; Nguyen and Csallner, 2015; Moran et al., 2018a), not opensourced (Xu et al., 2021), do not take website screenshot as input or do not produce complete website code (Asiroglu et al., 2019; Beltramelli, 2018). The details of prompting strategies used in our experiment as baselines are all available in our artifact.
@@ 6.2.1. High-Level Similarity
For code similarity, we adopt the Bilingual Evaluation Understudy (BLEU) score (Papineni et al., 2002) to evaluate the closeness of generated code $C_{g}$ to the original code $C_{0}$. BLEU score is an automatic evaluation metric widely used in machine translation studies. It calculates the similarity of machine-generated translations and human-created references (i.e., ground truth). It is also widely used in code generation tasks to measure the similarity between generated code and the ground truth  (Chen et al., 2021; Yin et al., 2018; Wei et al., 2019; Ren et al., 2020; Lu et al., 2021).
For image similarity, we measure image similarity between $I_{0}$ and $I_{g}$ via CLIP Score  (Si et al., 2024). CLIP score relies on the similarity of their CLIP (Radford et al., 2021) embeddings, denoted as CLIP($I_{0}$, $I_{g}$). Specifically, we extract features using CLIP-ViT-B/32 after resizing the screenshots to squares.
@@ 6.2.2. Fine-Grained Measurements
High-level metrics capture only the overall similarity of images and code, lacking the detail necessary to analyze specific model performances. To address this limitation, we adopt a suite of element-matching metrics for evaluating generated webpages in terms of text content, position, and color introduced by Si et al. (Si et al., 2024). Given reference and generated webpage screenshots $I_{0}$ and $I_{g}$, the algorithm detects visual element blocks and matches them using the Jonker-Volgenant algorithm. It then evaluates the similarity of the matching blocks across several aspects: (1) Block-match measures the ratio of matched block sizes to all block sizes, penalizing missing or hallucinated blocks; (2) Text similarity calculates character overlap using the Sørensen-Dice coefficient; (3) Color similarity employs the CIEDE2000 formula to evaluate perceptual color differences; and (4) Position similarity assesses the alignment of block centers.